OpenStudy (anonymous):

Help solving inequality: 0< 1/x-3 + 2/x-9 How would I do that in steps, I believe they want me to find domain only. Please show me steps.

OpenStudy (anonymous):

this is really a drag. you are going to have to add these up . no avoiding it

OpenStudy (anonymous):

you get \[0<\frac{1}{x-3}+\frac{2}{x-9}\] \[0<\frac{x-9+2(x-3)}{(x-3)(x-9)}\] \[0<\frac{3x-15}{(x-3)(x-9)}\] \[0<\frac{3(x-5)}{(x-3)(x-9)}\]

OpenStudy (anonymous):

zeros are at 3, 5 and 9 so you have to check the following intervals: \[(-\infty, 3)\] \[(3,5)\] \[(5,9)\] and \[(9,\infty)\]

OpenStudy (anonymous):

if you check you will see it is positive on \[(3,5)\cup (9,\infty)\]

OpenStudy (anonymous):

You deserve 10 medals! Do you tutor on skype by any chance?